OpenStudy (anonymous):

tan4Θ+tan2Θ/1-tan4Θtan2Θ a. tan6Θ b. cot6Θ c. tan4Θ d.cot4Θ

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[\tan(\alpha + \beta)=\frac{ \tan(\alpha)+\tan(\beta) }{ 1-\tan(\alpha)\tan(\beta) }\]How about\[\alpha=4\theta\]and \[\beta=2\theta\]Then the solution is\[\tan(6\theta)\]